NPP Projection Admin 10/23



To do this evening/ was completed: Just fyi - before anyone has questions about plattering Bugonia and our regular projection schedule, Alec and Matthew and I are working on this to see what we need overlap for and will keep you abreast of any changes, but for now, it's really looking like it just be me shifting to a mid for that opening week. A lot of this depends on how well it does and how much we'd like to schedule, but I'd also like to honor the fact that y'all have stuff outside of work so we'll make sure we're covered and happy.  - Lillian

A1 10/31 playlists built--Bugonia does not yet have the proper masking/lens presets included heads up. 10/17 playlists removed from servers.

New posters up, outgoing posters down.

Nashville and Jaws prints arrived, in T1. Inspected the first two reels of Nashville, decent condition but lots of splices.

Playlists made for Indiana Jones, Bride of Frankenstein (incomplete, will need updates). Made DCP of all the performance pieces for Bride's preshow just in case it turns out to be the easiest way to handle the show.

Halloween ingested to LMS via hard drive--playlist built with all of the performance cues attached.

Presentation Issues:

Preshow Clips/Playlist: Creature from the Black Lagoon, Imaginationland, Indiana Jones 3 all tech'd and sorted for this weekend. Just a friendly reminder to build playlists on the LMS for automation. - Lillian

Trailer Updates:

Special Events/ AV set up: Mics are in T3 for tomorrow. - L

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): Ingested Blade. Added tracking to distribution for Spartan. Lillian

Night of the Living Dead file from MY came in, so I've got it downloaded and in DCP-o-matic. Sound is analyzing. Check the volume and start transcoding this tonight. Thanks! - L

Outgoing/Shipping (content): Nashville and Jaws arrived - Brought down Scooby Doo with labels. It went out this afternoon. - Lillian

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning: Greg from LC is coming in on Wednesday morning to check out our projectors in T6 (and maybe take a look at some spare parts in T2). - Lillian 

We do in fact have another 30SP in T2, which I brought down to T1. Plug is still working well in T2...

I'll change out that lamp Monday. I think Monday will be a better time. - Lillian

Cue sheets updated/checked (every shift): Yes. Added weekend info for private events. - Lil

Marquee updated?: Multiple outgoing titles from yesterday all down, put up Shelby Oaks, Frankenstein, Springsteen, Blue Moon to the front. Took down Scooby and The Warriors, moved up Casper, added Innocent Blood and Nashville to the 35m, side. - Lillian 

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will